IN 8 years time I will be twice the age I was 3 years ago. How old am I now?
How will my equation look like?

Answer by mananth(16946)

I am x years now
3 years ago I was x-3 years
8 years hence I will be x+ 8 years

x+8=2(x-3)
x+8=2x-6
x=14